The length of a rectangle is 1ft less than twice its width. The area is 55ft^2 find the perimeter.
Answer by checkley77(12844)

L=2W-1
L*W=55
(2W-1)W=55
2W^2-W-55=0
(2W-11)(W+5)=0
2W=11
W=11/2
W=5.5 FT. ANS.
L=2*5.5-1
L=11-1
L=10 FT. ANS.
PROOF:
10*5.5=55
55=55
THUS THE PERIMETERIS:
2L+2W=PERIMETER.
2*10+2*5.5=20+11=31 FT. IS THE PERIMETER.